[figure: 9.7 cm]

SPONTANEOUS SURFACE MAGNETISATION OF MnF_{2} SINGLE CRYSTAL IN THE ANTIFERROMAGNETICSTATE

V. I. Nizhankovskii^{a}, A. I. Khar’kovskii^{a} and A. J. Zaleski^{b} aInternational Laboratory of High Magnetic Fields and Low Temperatures, Gajowicka 95 53-421 Wrocław, Poland bInstitute of Low Temperature and Structural Research, Polish Academy of Sciences, P.O. Box 937, 50-950 Wrocław, Poland

Spontaneous magnetisation of (100) and (010) surfaces of MnF_{2} single crystal in antiferromagnetic state is discovered. Its temperature dependence resembles that of a magnetic sublattice but the critical exponent of spontaneous magnetisation β_{S} is in the range 0.42 − 0.47 that is significantly different from the critical exponent for the sublattice magnetisation β = 0.31 obtained by the magnetic x-ray scattering measurements. Sign of the surface magnetisation is determined by difference in dielectric constants ϵ of MnF_{2} and ambient matter: magnetisation is always directed to the substance with smaller ϵ. Our explanation of above results is based on a simple microscopic model in which the influence of gradient of ϵ is reduced to the effective perturbation potential. The resulting decrease of the energy of the surface ions is compared with the energy of magnetic anisotropy. The lack of the spin-orbit interaction in the case of Mn^{2+} is favorable for the appearance of the surface magnetisation even for small gradients of the effective dielectric constant.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : V. I. Nizhankovskii

Address for correspondence : International Laboratory of High Magnetic Fields and Low Temperatures 53-421 Wroclaw, Gajowicka 95, Poland

Email address : nizhan@alpha.mlspmint.pan.wroc.pl